OWL shift summary for Dec 27, 1999

Richard Kessler (rkessler@fnal.gov)
Mon, 27 Dec 1999 07:07:50 -0600 (CST)

Shift Summary for OWL shift, Dec 27, 1999
========================================================================
On Shift: R.Kessler, U,Erkelslan
========================================================================
Hours of beam available at all intensities: 6.2 hr
Hours of beam available at nominal intensity: 6.2 hr
Hours of E832 data (special runs not included): 5.3 hr
Hours of special runs (laser scans, muon runs...): --
========================================================================
Accounting of lost time:
- 25 minutes due to piipeline cr 07 problems.
- 25 minutes for a few kfrends
========================================================================
Runs ended this shift:
Run 15304 - BEAM_799 - in progress

========================================================================
Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: None
========================================================================
Overview of shift:
E799 physics data at 9E12.

========================================================================
Description of Shift:

Great beam all shift ... but ...

At 2:40am the pipeline temp alarms appeared on every rack but everything
seemed OK and the happy-face thermometer at ETOT read 72 F. BobT thinks
it's a false alarm and recommends to ignore it and continue. About ten
minutes later we get PED EXP > 0 for CSI 947. Init pipeline hangs on
cr07, so we fart around for almost 1/2 hour to re-boot cr07. After
finally taking data again, csi 947 continues to have pedexp>0 and Val
continues to $#&^&$#^$. At about 3:30a.m. our sleepy Spokesman makes
an on-the-fly decision to continue with the bad channel, and to wait
for a 1 hour down-time before attempting to replace this channel.

MISC INFO about CSI 947 PEDEXP>0 :
- the pedexp>0 occured during Run 15304 at spill 154-.
- it is a small block near the big/small boundary.
- etot#sum macro is messed up due to LARGE [false] energy in csi 947.
- be aware that the CSI is used in L3, so simply ignoring
csi 947 off-line may not be sufficient.

TRIGGER LATCH CRATE 8:
I suspected that perhaps the busy signals were not treated
correctly in this crate. I injected a DC NIM level into both
DYC BUSY and DYC FULL and the trigger behaved correctly;
stopped in error state for DYC-BUSY, and just paused for DYC-FULL.
So still no clue about this crate. The test was done off-spill
so that no beam was lost.
Next step is to check input control signals: L1, L2, RF, etc ...

========================================================================
Follow up Notes on Previous Shift:
- No trouble from banana land, except for a few errors
in the kumac requesting e-mail [but NOT page] for expert.

========================================================================
Items to Follow up on Next Shift:
Swap DPMT 947 during >1 hour down-time.

========================================================================